Mogul Company ships merchandise to Ski Outfit in a consignment arrangement. The arrangement specifies that Ski Outfit will attempt to sell the merchandise, and in return, Mogul will pay to Ski Outfit a commission of 20% of the selling price on any merchandise sold. During the year, Mogul ships inventory with a cost of $92,000 to Ski Outfit and pays shipping costs of $16,400. By the end of the year, $72,000 of the merchandise has been sold to customers for a total of $97,000. Mogul allocates $12,000 of the shipping costs to inventory sold and the other $4,400 to inventory not sold. Mogul also paid advertising costs during the year of $16,000. What amount of inventory will Mogul report at year end?

Multiple Choice

  • A $44,400.

  • B $20,000.

  • C $60,400.

  • D $24,400.
Answer #1

cost of goods = 92000+16400=108400

cost of goods s old = 72000+12000=84000

inventory =108400-84000

=24400$

commission is selling expense and will not form part of invenoty.